# Learning Visuotactile Skills with Two Multifingered Hands

### v1

@@ II Related Work
Bimanual teleoperation. Having access to a diverse set of high-quality demonstrations is crucial for learning a high-quality policy. Existing teleoperation systems are largely restricted to the use of parallel-jaw grippers [1, 2, 27, 28] or single-handed scenarios [29, 30, 31]. In addition, the hand teleoperation systems heavily rely on retargeting to solve morphology differences, thus introducing unavoidable latency and are not intuitive to use. In contrast, our system teleoperates a robot hand by separating finger control into thumb control and power grasp control, providing a smoother user experience.
Learning visuotactile skills. Our system also features rich visuotactile sensory data. Integrated visuotactile sensing has been used for numerous applications such as grasping [32], in-hand manipulation [33], object shape reconstruction [34, 35], and cloth manipulation [36]. However, previous bimanual manipulation systems are either not equipped with dexterous hands or lack rich sensory signals. To the best of our knowledge, our work is the first at the intersection of bimanual dexterous manipulation and imitation learning from visuotactile inputs.
@@ V Experiments
Modality Default Init. Rare Init. Ours 10 / 10 10 / 10 w.o. Touch 10 / 10 4 / 10 w.o. Vision 10 / 10 0 / 10 TABLE II: Success rate on Tower Block Stacking task. We use two task initialization schemes (as illustrated in Figure 6). The default initialization (Default Init.) is more frequently encountered in the demonstration dataset compared to the other initialization (Rare Init.). Without visual or touch feedback, the policy struggles to succeed from Rare Init.     Modality Pickup Success Ours 10 / 10 5 / 10 w.o. Touch 10 / 10 0 / 10 w.o. Vision 0 / 10 0 / 10 EEF Only 0 / 10 0 / 10 TABLE III: Success rate on Steak Serving task using policies trained with different sensing modalities. The pickup success is an intermediate metric that measures how often the hands successfully pick up both objects. The policy cannot finish the task without all three sensing modalities and cannot pick up the object without touch.     Modality Pickup Success Ours 10 / 10 10 / 10 w. Depth 10 / 10 1 / 10 Only Wrist 0 / 10 1 / 10 Only 3rd View 0 / 10 0 / 10 TABLE IV: Success rate on Steak Serving task with different camera configurations. The pickup success is an intermediate metric that measures how often the hands successfully pickup both objects. We find having depth does not improve the performance and having all the cameras are necessary to finish this challenging task.
@@ V-D Importance of Vision and Touch
Vision and touch improve policy robustness. To further understand how the sensing modality affects the robustness of our policy, we experiment with a less common scene initialization for the Tower Block Stacking. Specifically, we rotate the blocks in random directions. Figure 6 shows a comparison between the default initialization and the rare initialization. This scene initialization is less encountered in the demonstration dataset, and the perturbed block configuration makes the two-block pile harder to be picked up by two robot hands. In Table IV, we show a comparison of the success rate for this rare initialization and the default initialization across three different sensing configurations. While touch and vision are not needed for the default scene initialization (they all achieve 100% success rate), for the rare scene initialization, the policy trained without touch can only succeed 4/10 times, and the policy trained without vision cannot succeed at all. This suggests that vision and touch sensing modalities allow the policy to be more robust to rare scenarios.
@@ VI Discussions
Figure 6: Two scene initializations for the Tower Block Stacking task. The default initialization is shown on the left and the rare initialization is shown on the right. In the rare initialization, the perturbed block configuration makes the two-block pile harder to pick up by two robot hands and more difficult to stack stably on another block.

### v2

@@ II Related Work
Bimanual teleoperation. Having access to a diverse set of high-quality demonstrations is crucial for learning a high-quality policy. Existing teleoperation systems are largely restricted to the use of parallel-jaw grippers [1, 2, 27, 28, 29, 30] or single-handed scenarios [31, 32, 33, 34]. In addition, the hand teleoperation systems heavily rely on retargeting to solve morphology differences, thus introducing unavoidable latency and are not intuitive to use. In contrast, our system teleoperates a robot hand by separating finger control into thumb control and power grasp control, providing a smoother user experience.
Learning visuotactile skills. Our system also features rich visuotactile sensory data. Integrated visuotactile sensing has been used for numerous applications such as grasping [35], in-hand manipulation [36], object shape reconstruction [37, 38, 39] and object recognition [40], cloth manipulation [41], and learning representation for object interactions [42, 43]. However, none of them use two multifingered hands, and previous bimanual manipulation systems are either not equipped with dexterous hands or lack rich sensory signals. To the best of our knowledge, our work is the first at the intersection of bimanual dexterous manipulation and imitation learning from visuotactile inputs.
@@ V Experiments
Modality Default Init. Rare Init. Ours 10 / 10 10 / 10 w.o. Touch 10 / 10 4 / 10 w.o. Vision 10 / 10 0 / 10 TABLE II: Success rate on Block Stacking task. We use two task initialization schemes (as illustrated in Figure 6). The default initialization (Default Init.) is more frequently encountered in the demonstration dataset compared to the other initialization (Rare Init.). Without visual or touch feedback, the policy struggles to succeed from Rare Init.     Modality Pickup Success Ours 10 / 10 5 / 10 w.o. Touch 10 / 10 0 / 10 w.o. Vision 0 / 10 0 / 10 EEF Only 0 / 10 0 / 10 TABLE III: Success rate on Steak Serving task using policies trained with different sensing modalities. The pickup success is an intermediate metric that measures how often the hands successfully pick up both objects. The policy cannot finish the task without all three sensing modalities and cannot pick up the object without touch.     Modality Pickup Success Ours 10 / 10 10 / 10 w. Depth 10 / 10 1 / 10 Only Wrist 0 / 10 1 / 10 Only 3rd View 0 / 10 0 / 10 TABLE IV: Success rate on Steak Serving task with different camera configurations. The pickup success is an intermediate metric that measures how often the hands successfully pickup both objects. We find having depth does not improve the performance and having all the cameras are necessary to finish this challenging task.
@@ V-D Importance of Vision and Touch
Vision and touch improve policy robustness. To further understand how the sensing modality affects the robustness of our policy, we experiment with a less common scene initialization for the Block Stacking. Specifically, we rotate the blocks in random directions. Figure 6 shows a comparison between the default initialization and the rare initialization. This scene initialization is less encountered in the demonstration dataset, and the perturbed block configuration makes the two-block pile harder to be picked up by two robot hands. In Table IV, we show a comparison of the success rate for this rare initialization and the default initialization across three different sensing configurations. While touch and vision are not needed for the default scene initialization (they all achieve 100% success rate), for the rare scene initialization, the policy trained without touch can only succeed 4/10 times, and the policy trained without vision cannot succeed at all. This suggests that vision and touch sensing modalities allow the policy to be more robust to rare scenarios.
@@ VI Discussions
Figure 6: Two scene initializations for the Block Stacking task. The default initialization is shown on the left and the rare initialization is shown on the right. In the rare initialization, the perturbed block configuration makes the two-block pile harder to pick up by two robot hands and more difficult to stack stably on another block.